OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (14, -2) and is perpendicular to 7x – 10y = 18

OpenStudy (anonymous):

Slope intercept is y=mx + b If you can first change your 7x - 10y = 18 equation to follow the format y=mx +b, you can find out your slope (m). For a line to be perpendicular to another line, it must be the negative reciprocal..

OpenStudy (anonymous):

7x - 10y = 18 -10y = -7x + 18 y = (7/10)x + 18 and thus the slope for the 7x - 10y = 18 equation is 7/10. To find the slope of the perpendicular equation you are looking for, you take the negative reciprocal = -10/7

OpenStudy (anonymous):

That will equal m. Then plug in the points and other info you know to the formula y-y1=m(x-x1) and simplify.

OpenStudy (anonymous):

y = mx + c for perpendicular lines: \[ m_1 m_2 = -1 \] the given equation : \[ y = 0.7x - 1.8 \] The equation u want must be perpendicular to this \[ \Rightarrow m = \frac{-1}{0.7} = -1.428 \] \[ \Rightarrow y = c_2 - 1.428 x \] But this satisfies y = -2 for x = 14 \[ -2 = c_2 - 20 \Rightarrow c_2 = -22 \Rightarrow y = -1.428 x -22 \Rightarrow 7y + 10x + 154 = 0 \]